Determine whether a system of two linear equations with (0,0) and (2,2) as solutions sometimes, always, or never has other solutions.
Answer by oberobic(2304)

A line is defined by two points. You have two points, so you have one line.
.
We know two points: (0,0) and (2,2)
m = slope = rise/run = change in y/ change in x
m = (2-0)/(2-0) = 2/2 = 1
.
By inspection we know the x and y intercepts, which is the origin (0,0).
.
That means the equation is:
y = mx +b
y = 1*x +0
y = x
.